Two industrial cleaning products salesman A and B, must each make two calls per day, one in the morning and one in the afternoon. Salesman A has a probability of 0.45 of selling a cleaner on any call, while salesman B (a novice) has a probability of 0.2 of a sale. The two salesman work independently of each other, and each salesman’s morning and afternoon results are independent of each other.
2. *Question*:
Find the probability that salesman B makes one sale for the day (rounded off to two decimals)?

Answer by ikleyn(53765)

.

P = P(one sale in the morning or one sale in the afternoon) - P(both) = 0.2 + 0.2 - 0.2*0.2 = 0.4 - 0.04 = 0.36.